# Euclid Quick Data Release (Q1) point-like X-ray selected sources

### v1

@@
To better understand the role of active galactic nuclei (AGN) in galaxy evolution, it is crucial to achieve a complete and pure AGN census. X-ray surveys are key to this, but identifying their counterparts (CTPs) at other wavelengths remains challenging due to their larger positional uncertainties and limited availability of deeper, uniform ancillary data. Euclid is revolutionising this effort, offering extensive coverage of nearly the entire extragalactic sky, particularly in the near-infrared bands, where AGN are more easily detected. With the first Euclid Quick Data Release (Q1), we tested and validated methods for identifying, classifying, and determining the redshifts of Euclid CTPs to known point-like sources from major X-ray surveys, including XMM-Newton, Chandra, and eROSITA. Using Bayesian statistics, combined with machine learning (ML), as incorporated in the algorithm NWAY, we identify the CTPs to 11 286 X-ray sources from the three X-ray telescopes. For the large majority of 10 194 sources, the association is unique, with the remaining 10% of multi-CTP cases equally split between XMM-Newton and eROSITA. Six percent of the Euclid CTPs are detected in more than one X-ray survey. ML is then used to distinguish between Galactic (8%) and extragalactic (92%) sources. We computed photo-s using deep learning for the 8617 sources detected in the 10th data release of the DESI Legacy Survey, reaching an accuracy and a fraction of outliers of about 5%. Based on their X-ray luminosities, over 99% of CTPs identified as extragalactic are classified as AGN, most of which appear unobscured according to their hardness ratios. With this paper, we release our catalogue, which includes identifiers, basic X-ray properties, the reliability of the associations, and additional property extensions such as Galactic/extragalactic classifications and photometric/spectroscopic redshifts. We also provide probabilities for sub-selecting the sample based on purity and completeness, allowing users to tailor the sample according to their specific needs.
@@
Active galactic nuclei (AGN111Used interchangeably in both singular and plural cases), which rank among the most energetic phenomena in the Universe, play a pivotal role in shaping the evolutionary trajectories of galaxies throughout cosmic time. Their activity is fundamentally driven by the accretion of matter onto supermassive black holes (SMBHs) from their surrounding environment (e.g., Salpeter 1964; Lynden-Bell 1969; Pringle & Rees 1972; Shakura & Sunyaev 1973; Lynden-Bell & Pringle 1974; Kormendy & Ho 2013). Situated at the centres of galaxies, AGN emit intense radiation across the entire electromagnetic spectrum, ranging from radio waves to -rays (e.g., Elvis et al. 1994; Urry & Padovani 1995; Padovani et al. 2017). While this is particularly true for quasi-stellar objects (QSOs), it is important to note that the range of activity, emission, and interaction of AGN spans a broad spectrum, with different AGN exhibiting varying levels of energy outputs (e.g., Peterson 1997; Pierce et al. 2010; Pović et al. 2012; Bettoni et al. 2015).
@@
Consequently, with enough coupling efficiency, this allows AGN to exert significant influence over both their host galaxies and the growth of their central SMBHs, where the interplay between AGN activity and the respective host is believed to regulate star formation and shape the morphology of galaxies. Here, energetic feedback drives powerful winds and outflows, impacting the surrounding interstellar and intergalactic medium (Greene et al. 2011; Cielo et al. 2018). These close connections are supported by fundamental scaling relations that link the SMBH mass to various galactic properties. Notable among these are the – relation, which ties BH mass to the velocity dispersion of stars in the galaxy’s bulge (Gebhardt et al. 2000; Ferrarese & Merritt 2000), and the – relation, which connects it to the stellar mass of the host galaxy (Magorrian et al. 1998; Häring & Rix 2004; Gültekin et al. 2009; Sani et al. 2011; Shankar et al. 2016; Suh et al. 2020), pointing to an intricate co-evolution of the systems (Kormendy & Ho 2013; Heckman & Best 2014; Madau & Dickinson 2014; Euclid Collaboration: Bisigello et al. 2024). However, the physical mechanisms driving this connection remain under debate, highlighting the need to construct a complete picture of AGN diversity and complexity to better understand galaxy formation and transformation across the history of the Universe (e.g., Aird et al. 2010; Buchner et al. 2015; Georgakakis et al. 2015; Ananna et al. 2017; Morganti 2017; Delvecchio et al. 2017; Husemann & Harrison 2018; Harrison & Ramos Almeida 2024).
@@
Despite significant advancements over the years, the current AGN census remains incomplete due to inherent biases introduced to individual samples by various multi-wavelength selection methods, each favouring specific properties of the population (Messias et al. 2014; Padovani et al. 2017; Delvecchio et al. 2017; Lyu et al. 2022; Green et al. 2024). Comprehensive wavelength coverage is essential for identifying diverse AGN types across all redshifts, with each method capitalising on the dominant physical processes within its waveband to uncover unique facets of AGN activity, where the AGN dominates the spectral energy distribution (SED) of its host galaxy (Padovani et al. 2017). For instance, radio observations are effective in identifying jet-dominated AGN (De Breuck et al. 2002; Smolčić et al. 2017), while optical diagnostics (Feltre et al. 2016), such as emission line ratios (e.g., BPT diagrams as in Baldwin et al. 1981; Veilleux & Osterbrock 1987; Juneau et al. 2014), are widely used to distinguish AGN from star-forming galaxies by probing ionisation mechanisms. Other selections include colour or variability (e.g., Richards et al. 2002; Bongiorno et al. 2010; Bovy et al. 2012; Donley et al. 2012; Kirkpatrick et al. 2013; Peters et al. 2015; Palanque-Delabrouille et al. 2016; Lusso & Risaliti 2016). Soft X-ray selection excels at detecting unobscured accretion-powered emission (e.g., Hasinger 2008; Brandt & Hasinger 2005; Nandra et al. 2015) and can be complemented by mid-infrared (MIR) observations (e.g., Stern et al. 2012; Assef et al. 2013), which penetrate dense dust clouds effectively (Hickox & Alexander 2018). Notably, the significantly higher contrast between the intrinsic brightness of accreting SMBHs and their host galaxies in X-rays, as opposed to other wavelengths, makes this selection one of the most uncontaminated, providing pure AGN samples of larger diversity (Xue et al. 2011a; Donley et al. 2012). To combine different observational perspectives of AGN, these X-ray samples must then be paired with ancillary multi-wavelength data.
@@
In the past, identifying multi-wavelength counterparts (CTPs) for X-ray selected sources was a significant hurdle in AGN characterisation and redshift estimation (Salvato et al. 2018). While combining data from multiple wavelengths is crucial for constructing a more complete picture of AGN, the process of cross-identifying sources remains complicated. Due to the significant positional uncertainties in X-ray surveys, confidently linking most X-ray sources to a single, definitive CTP in e.g., optical, infrared, or radio surveys is challenging, except for the brightest cases. These uncertainties, along with the uneven spatial coverage of ancillary data, further complicate the task of reliably pinpointing AGN across different surveys. Consequently, a simple coordinate match is insufficient, and more sophisticated methods are needed for accurate cross-identification. For this purpose, machine learning (ML) algorithms that integrate multi-wavelength source classification approaches have become increasingly prevalent (e.g., Cavuoti et al. 2014; Brescia et al. 2015; Karsten et al. 2023; Cooper et al. 2023; Daoutis et al. 2023; Zeraatgari et al. 2024; Pérez-Díaz et al. 2024; Mechbal et al. 2024).
@@
While challenging to this date, the European Space Agency’s Euclid satellite will help overcome this limitation due to its uniform depth and wide coverage (Laureijs et al. 2011; Euclid Collaboration: Mellier et al. 2024). At the same time, Euclid’s depth results in very high source densities, increasing the risk of unforced misidentification and chance alignments. Over its six-year mission lifetime, Euclid will cover a total of 14 000 deg2 in the Euclid Wide Survey (EWS, Euclid Collaboration: Scaramella et al. 2022), expected to detect billions of sources. In addition to its cosmological objectives, Euclid will play a crucial role in the detection and characterization of AGN, expecting to detect at least 10 million of them (Euclid Collaboration: Bisigello et al. 2024; Euclid Collaboration: Lusso et al. 2024; Euclid Collaboration: Selwood et al. 2025). To achieve these goals, Euclid employs two main instruments: the VISible imager (VIS, Euclid Collaboration: Desprez et al. 2020; Euclid Collaboration: Cropper et al. 2024), covering wavelengths from 0.53 to 0.90 () and the Near-Infrared Spectrometer and Photometer (NISP, Maciaszek et al. 2022; Euclid Collaboration: Jahnke et al. 2024) operating in the near-infrared, with photometric measurements in three bands (, , and ), covering in total the wavelength range from 0.95 to 2.02 , while also including a slitless spectrograph for the detection of emission lines.
@@
With the first Euclid Quick Data Release (Q1) we can now begin building and testing the necessary machinery for identifying and characterising AGN previously selected in the X-rays, where we aim to leverage Euclid’s depth, resolution and wavelength coverage to provide more secure associations. This initial testing will form the foundation for refining our methods and ensuring accurate AGN identification as more data becomes available over the next years.
@@
In this paper, we address the task of reliably associating X-ray sources with their Euclid CTPs and investigate their respective NIR properties as a function of X-ray flux, among other properties. The paper is structured as follows: Sect. 2 gives an overview of the Q1 data. Section 3 introduces the X-ray source catalogues used in this work, while Sect. 4 outlines the approach to identifying CTPs. Section 5 describes how using a sample of sources clearly characterised with the 10th Data Release of the Legacy Survey (Dey et al. 2019), we assign a Galactic/extragalactic nature to our sources using Euclid photometry. In Sect. 6, we compute photometric redshifts for all CTPs. Section 7 displays CTP sample properties and Sect. 8 explains the release of the catalogues and how to use them. A summary and discussion are given in the final Sect. 9.
@@
In this paper, unless stated otherwise, we express magnitudes in the AB system (Oke & Gunn 1983) and adopt a flat CDM cosmology with km s-1 Mpc-1, and , to facilitate direct comparison with similar works from literature.
@@ 2 Euclid Q1 fields
Euclid Q1 marks the initial data release with a footprint area of 63.1 deg2 (Euclid Collaboration: Mellier et al. 2024). It consists of observations making up the Euclid Deep Survey (EDS), selected for their considerable multi-wavelength coverage. However, the data is currently at the depth of the EWS, with a 5 depth of (AB)24.5 for point sources. The EDS targets three important Euclid Deep Fields (EDFs, see Fig. 1): the EDF Fornax (EDF-F), EDF South (EDF-S), and the EDF North (EDF-N). For more details, we refer the reader to Euclid Collaboration: Aussel et al. (2025).
@@ 3 The X-ray surveys
X-ray surveys are often constrained by the flux detection limit of the most sensitive X-ray telescopes, which typically operate with small fields of view. Therefore, surveys with different observational strategies need to be combined to facilitate the characterisation of the full AGN population. Pencil-beam surveys that provide exceptionally deep observations, valuable for probing faint AGN and for characterising the low-luminosity end of the AGN population, are limited in sky coverage and, consequently, miss rare (bright or distant) sources (e.g., Brusa et al. 2010; Hsu et al. 2014; Nandra et al. 2015; Marchesi et al. 2016; Luo et al. 2017; Oh et al. 2018). Subsequently, the data releases of XMM-Newton 4XMM DR14 (Webb et al. 2020) and Chandra Source Catalogue 2 (CSC 2.0, Evans et al. 2024) feature outstanding sensitivity for X-ray selected sources but cover relatively small regions of the sky.
On the other hand, wide-area surveys are crucial for complementing pointed observations and uncover those AGN missed (e.g., Ananna et al. 2017; Fotopoulou et al. 2016). Therefore, we incorporate data from the extended Roentgen Survey with an Imaging Array (eROSITA; Predehl et al. 2021) DR1 Main sample (Merloni et al. 2024). As part of the ongoing eROSITA All-Sky Survey (eRASS), these data provide homogeneous, soft X-ray coverage of the entire extragalactic sky. While eROSITA’s positional accuracy is somewhat pourer than that of XMM-Newton and Chandra, its unparalleled breadth allows us to include a significantly larger sample of sources, enhancing the statistical power of our analysis and ensuring a more comprehensive AGN census (see Fig. 1 and Footnote 2). In the following subsections, we describe the selection criteria applied to the catalogues from each X-ray survey.
X-ray catalogue | Number of sources | AND point-like | AND POSERR ¡ limit | EDF-F | EDF-N | EDF-S |
(1) | (2) | (3) | (4) | (5) | (6) | (7) |
4XMM DR14 | 692 109 | 612 521 | 612 504 | 4523 | 445 | 273 |
CSC 2.0 | 317 167 | 296 473 | 296 162 | 1016 | 506 | 147 |
eROSITA DR1 main | 898 812 | 879 153 | 878 984 | 1012 | 0 | 3369 |

### v2

@@
To better understand the role of active galactic nuclei (AGN) in galaxy evolution, it is crucial to work with a complete and pure AGN sample. X-ray surveys are key to doing so, but their larger positional uncertainties complicate counterpart (CTP) association, further compounded by the limited availability of deep, uniform multi-wavelength ancillary data. Euclid is revolutionising this identification effort, offering extensive coverage of nearly the entire extragalactic sky, particularly in the near-infrared bands, where AGN are more easily detected. Using the first Euclid Quick Data Release (Q1), we validated the methods for identifying and classifying Euclid CTPs of known point-like sources from major X-ray surveys, including XMM-Newton, Chandra, and eROSITA. Using Bayesian statistics, combined with machine learning (ML), as incorporated in the algorithm NWAY, we identified the CTPs of 11 286 X-ray sources from the three X-ray telescopes. For the large majority of 10 194 sources, the association is unique, with the remaining 10% of multi-CTP cases equally split between XMM-Newton and eROSITA. Six percent of the Euclid CTPs are detected in more than one X-ray survey. We then used ML to distinguish between Galactic (8%) and extragalactic (92%) sources. We computed photo-s using deep learning for the 9259 sources detected in the tenth data release of the DESI Legacy Survey, reaching an accuracy and a fraction of outliers of roughly 5%. Based on their X-ray luminosities, all CTPs identified as extragalactic are classified as AGN, most of which appear as type I AGN according to their hardness ratios. With this paper, we release our catalogue, which includes identifiers, basic X-ray properties, the reliability of the associations, and additional property extensions, such as Galactic- or extragalactic classifications and photometric/spectroscopic redshifts. We also provide probabilities for sub-selecting the sample based on purity and completeness, in order to allow users to tailor the sample according to their specific needs.
@@
Active galactic nuclei (AGN111We use this term interchangeably to indicate both the singular and plural cases.), which rank among the most energetic phenomena in the Universe, play a pivotal role in shaping the evolutionary trajectories of galaxies throughout cosmic time. Their activity is fundamentally driven by the accretion of matter onto supermassive black holes (SMBHs) from their surrounding environment (e.g., Salpeter 1964; Lynden-Bell 1969; Pringle & Rees 1972; Shakura & Sunyaev 1973; Lynden-Bell & Pringle 1974; Kormendy & Ho 2013). Situated at the centres of galaxies, AGN emit intense radiation across the entire electromagnetic spectrum, ranging from radio waves to -rays (e.g., Elvis et al. 1994; Urry & Padovani 1995; Padovani et al. 2017). While this is particularly true for quasi-stellar objects (QSOs), it is important to note that the range of activity, emission, and interaction of AGN span a broad spectrum, with different AGN exhibiting varying levels of energy outputs (e.g., Peterson 1997; Pierce et al. 2010; Pović et al. 2012; Bettoni et al. 2015).
@@
Consequently, with enough coupling efficiency, this allows AGN to exert significant influence over both their host galaxies and the growth of their central SMBHs, where the interplay between AGN activity and the respective host is believed to regulate star formation and shape the morphology of galaxies. Here, energetic feedback drives powerful winds and outflows, impacting the surrounding interstellar and intergalactic medium (Greene et al. 2011; Cielo et al. 2018). These close connections are supported by fundamental scaling relations that link the SMBH mass to various galactic properties. Notable among these are the - relation, which ties black hole (BH) mass to the velocity dispersion of stars in the galaxy’s bulge (Gebhardt et al. 2000; Ferrarese & Merritt 2000), and the - relation, which connects it to the stellar mass of the host galaxy (Magorrian et al. 1998; Häring & Rix 2004; Gültekin et al. 2009; Sani et al. 2011; Shankar et al. 2016; Suh et al. 2020), thus pointing to an intricate co-evolution of the systems (Kormendy & Ho 2013; Heckman & Best 2014; Madau & Dickinson 2014; Euclid Collaboration: Bisigello et al. 2024). However, the physical mechanisms driving this connection remain under debate, highlighting the need to construct a complete picture of AGN diversity and complexity to better understand galaxy formation and transformation across the history of the Universe (e.g., Aird et al. 2010; Buchner et al. 2015; Georgakakis et al. 2015; Ananna et al. 2017; Morganti 2017; Delvecchio et al. 2017; Husemann & Harrison 2018; Harrison & Ramos Almeida 2024).
@@
Despite significant advancements over the years, the current AGN census remains incomplete due to inherent biases introduced to individual samples by various multi-wavelength selection methods, as each favours specific properties of the population (Messias et al. 2014; Padovani et al. 2017; Delvecchio et al. 2017; Lyu et al. 2022; Green et al. 2024). Comprehensive wavelength coverage is vital for identifying diverse AGN across redshift, as each method that contributes to this coverage capitalises on the dominant physical processes of AGN activity, dominating the spectral energy distribution (SED) of its host galaxy (Padovani et al. 2017). For instance, radio observations are effective at identifying jet-dominated AGN (De Breuck et al. 2002; Smolčić et al. 2017), while optical diagnostics (Feltre et al. 2016), such as emission line ratios (e.g., BPT diagrams as in Baldwin et al. 1981; Veilleux & Osterbrock 1987; Juneau et al. 2014), are widely used to distinguish AGN from star-forming galaxies by probing ionisation mechanisms. Other selections include colour or variability (e.g., Richards et al. 2002; Bongiorno et al. 2010; Bovy et al. 2012; Donley et al. 2012; Kirkpatrick et al. 2013; Peters et al. 2015; Palanque-Delabrouille et al. 2016; Lusso & Risaliti 2016). Soft X-ray selection excels at detecting unobscured accretion-powered emission (e.g., Hasinger 2008; Brandt & Hasinger 2005; Nandra et al. 2015) and can be complemented by mid-infrared (MIR) observations (e.g., Stern et al. 2012; Assef et al. 2013), which can effectively penetrate dense dust clouds (Hickox & Alexander 2018). Notably, the significantly higher contrast between the intrinsic brightness of accreting SMBHs and their host galaxies in X-rays, as opposed to other wavelengths, makes this selection one of the most uncontaminated, and thus it provides pure AGN samples of a larger diversity (Xue et al. 2011a; Donley et al. 2012). To combine different observational perspectives of AGN, these X-ray samples must then be paired with ancillary multi-wavelength data.
@@
In the past, identifying multi-wavelength counterparts (CTPs) for X-ray-selected sources was a significant hurdle in AGN characterisation and redshift estimation (Salvato et al. 2018). While combining data from multiple wavelengths is crucial for constructing a more complete picture of AGN, the process of cross-identifying sources remains complicated. Due to the significant positional uncertainties in X-ray surveys, confidently linking most X-ray sources to a single, definitive CTP in, for example, optical, infrared, or radio surveys, is challenging, except in the brightest cases. These uncertainties, along with the uneven spatial coverage of ancillary data, further complicate the task of reliably pinpointing AGN across different surveys. Consequently, a simple coordinate match is insufficient, and more sophisticated methods are needed for accurate cross-identification. For this purpose, machine learning (ML) algorithms that integrate multi-wavelength source classification approaches have become increasingly prevalent (e.g., Cavuoti et al. 2014; Brescia et al. 2015; Karsten et al. 2023; Cooper et al. 2023; Daoutis et al. 2023; Zeraatgari et al. 2024; Pérez-Díaz et al. 2024; Mechbal et al. 2024).
@@
The European Space Agency’s Euclid satellite will also help overcome this limitation due to its uniform depth and wide coverage (Laureijs et al. 2011; Euclid Collaboration: Mellier et al. 2025). At the same time, Euclid’s depth results in very high source densities, thus increasing the risk of unforced misidentification and chance alignments. Over its six-year mission lifetime, Euclid will cover a total of 14 000 deg2 in the Euclid Wide Survey (EWS, Euclid Collaboration: Scaramella et al. 2022), and it is expected to detect billions of sources. In addition to its cosmological objectives, Euclid will play a crucial role in the detection and characterisation of AGN, as at least 10 million AGN are anticipated to be detected (Euclid Collaboration: Bisigello et al. 2024; Euclid Collaboration: Lusso et al. 2024; Euclid Collaboration: Selwood et al. 2025). To achieve these goals, Euclid employs two main instruments: the VISible imager (VIS, Euclid Collaboration: Desprez et al. 2020; Euclid Collaboration: Cropper et al. 2025), covering wavelengths from 0.53 to 0.90 (), and the Near-Infrared Spectrometer and Photometer (NISP, Maciaszek et al. 2022; Euclid Collaboration: Jahnke et al. 2025). Operating in the near-infrared and with photometric measurements in three bands (, , and ), Euclid covers in total the wavelength range from 0.95 to 2.02 while also including a slitless spectrograph for the detection of emission lines.
@@
With the first Euclid Quick Data Release (Q1), we can now begin building and testing the necessary machinery for identifying and characterising AGN previously selected in the X-rays. We aim to leverage Euclid’s depth, resolution, and wavelength coverage to provide more secure associations of these AGN. Moreover, this initial testing will form the foundation for refining our methods and ensuring accurate AGN identification as more data become available over the next years.
@@
In this paper, we address the task of reliably associating X-ray sources with their Euclid CTPs and investigate their respective NIR properties as a function of X-ray flux, among other properties. The paper is structured as follows: In Sect. 2 we give an overview of the Q1 data. In Sect. 3 we introduce the X-ray source catalogues used in this work, while Sect. 4 provides an outline of the approach to identifying CTPs. In Sect. 5, using a sample of sources clearly characterised with the tenth Data Release of the Legacy Survey (Dey et al. 2019), we describe how we assign a Galactic- or extragalactic nature to our sources using Euclid photometry. In Sect. 6, we compute photometric redshifts for all CTPs. Section 7 displays CTP sample properties, and in Sect. 8 we explain the release of the catalogues and how to use them. Finally, a summary and discussion are given in Sect. 9.
@@
In this paper, unless stated otherwise, we express magnitudes in the AB system (Oke & Gunn 1983) and adopt a flat CDM cosmology with km s-1 Mpc-1, , and to facilitate direct comparison with similar works from literature.
@@ 2 Euclid Q1 fields
Euclid Q1 marks the initial data release with a footprint area of 63.1 deg2 (Euclid Collaboration: Mellier et al. 2025). It consists of observations making up the Euclid Deep Survey (EDS), selected for their considerable multi-wavelength coverage. However, the data is currently at the depth of the EWS, with a 5 depth of (AB)24.5 for point sources. The EDS targets three important Euclid Deep Fields (EDFs; see Fig. 1): the EDF Fornax (EDF-F), EDF South (EDF-S), and the EDF North (EDF-N). For more details, we refer the reader to Euclid Collaboration: Aussel et al. (2025).
@@ 3 The X-ray surveys
X-ray surveys are often constrained by the flux detection limit of the most sensitive X-ray telescopes, which typically operate with small fields of view. Therefore, surveys with different observational strategies need to be combined to facilitate the characterisation of the full AGN population. Pencil-beam surveys that provide exceptionally deep observations, valuable for probing faint AGN and for characterising the low-luminosity end of the AGN population, are limited in sky coverage and, consequently, miss rare (bright or distant) sources (e.g., Brusa et al. 2010; Hsu et al. 2014; Nandra et al. 2015; Marchesi et al. 2016; Luo et al. 2017; Oh et al. 2018). Subsequently, the data releases of XMM-Newton 4XMM DR14 (Webb et al. 2020) and Chandra Source Catalogue 2 (CSC 2.0, Evans et al. 2024) feature outstanding sensitivity for X-ray-selected sources but cover relatively small regions of the sky.
On the other hand, wide-area surveys are crucial for complementing pointed observations and uncovering those AGN missed (e.g., Ananna et al. 2017; Fotopoulou et al. 2016). Therefore, we incorporate data from the extended Roentgen Survey with an Imaging Array (eROSITA; Predehl et al. 2021) DR1 main sample (Merloni et al. 2024). As part of the ongoing eROSITA All-Sky Survey (eRASS), these data provide homogeneous, soft X-ray coverage of the entire extragalactic sky. While eROSITA’s positional accuracy is somewhat poorer than that of XMM-Newton and Chandra, its unparalleled breadth allows us to include a significantly larger sample of sources, enhancing the statistical power of our analysis and ensuring a more comprehensive AGN census (see Fig. 1 and Footnote 2). In the following subsections, we describe the selection criteria applied to the catalogues from each X-ray survey.
| X-ray catalogue | MOC | Number of sources | AND point-like | AND POSERR ¡ limit | EDF-F | EDF-N | EDF-S |
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) |
| Unit | deg2 | ||||||
| 4XMM DR14 | 6.18 | 692 109 | 612 521 | 612 504 | 4523 | 445 | 273 |
| CSC 2.0 | 1.31 | 317 167 | 296 473 | 296 162 | 1016 | 506 | 147 |
| eROSITA DR1 main | 43.14 | 898 812 | 879 153 | 878 984 | 1012 | 0 | 3369 |